[figure: 9.7 cm]

QUANTUM TRANSFER MATRIX STUDY OF THE SPIN-PEIERLS COMPOUND CuGeO_{3}

M. Bieliński Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland e-mail: biel@pearl.amu.edu.pl

Thermodynamical properties of one-dimensional S=1/2 Heisenberg model with dimerized nearest and uniform next-nearest neighbour interactions are studied by the numerically exact quantum transfer matrix method and the results are applied to CuGeO_{3}. Suzuki-Trotter formula is used to obtain a classical system with spin σ=3/2 and effective interactions between nearest neigbours only. Magnetic susceptibility curve is calculated and compared with experimental results performed in a wide temperature range, revealing the presence of frustration in the model proposed for CuGeO_{3}. Temperature dependence of the dimerization parameter below the spin-Peierls transition point is also estimated.

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : M. Bieliński

Address for correspondence : Institute of Physics, Adam Mickiewicz University, ul. Umultowska 85, 61-614 Poznań, Poland

Email address : biel@pearl.amu.edu.pl